Determine the mass of 2.04% NaF solution that contains 3.67 g of F-.
Answer #1


Molar mass of F = 19 g/mol


mass(F)= 3.67 g

use:
number of mol of F,
n = mass of F/molar mass of F
=(3.67 g)/(19 g/mol)
= 0.1932 mol
This is number of moles of F

one mole of NaF contains 1 moles of F
use:
number of moles of NaF = number of moles of NaF / 1
= 0.1932 / 1
= 0.1932

Molar mass of NaF,
MM = 1*MM(Na) + 1*MM(F)
= 1*22.99 + 1*19.0
= 41.99 g/mol

use:
mass of NaF,
m = number of mol * molar mass
= 0.1932 mol * 41.99 g/mol
= 8.111 g

Now use:
Mass % of NaF = mass of NaF * 100 / mass of solution
2.04 = 8.111 * 100 / mass of solution
mass of solution = 398 g
Answer: 398 g